E3-5 Devin Wolf Company has the following balances in selected accounts on December 31,2017. Accounts Receivable $-0- Accumulated Depreciation-Equipment -0- Equipment 7,000 Interest Payable -0- Notes Payable 10,000 Prepaid Insurance 2,100 Salaries and Wages Payable -0- Supplies 2,450 Unearned Service Revenue 30,000 All the accounts have normal balances. The information below has been gathered at December 31,2017 1 . Devin Wolf Company borrowed $10,000 by signing a 9%, one-year note on September 1, 2017 2. A count of supplies on December 31, 2017, indicates that supplies of $900 are on hand 3. Depreciation on the equipment for 2017 is $1,000, 4. Devin Wolf Company paid $2,100 for 12 months of insurance coverage on June 1, 2017, 5. On December 1,2017, Devin Wolf collected $32,000 for consulting services to be performed from December 1, 2017, through March 31, 2018. 6. Devin Wolf performed consulting services for a client in December 2017. The client will be billed $4,200. 7. Devin Wolf Company pays its employees total salaries of $9,000 every Monday for the preceding 5-day week (Monday through Friday). On Monday, December 29, employees were paid for the week ending December 26. All employees worked the last 3 days of 2017 Instructions Prepare adjusting entries for the seven items described above
Answer #1

Solution:

1.Devin Wolf company borrowed$10,000 signing a 9% one year note on September 1,2017.

Debit Credit
Cash $10,000
Notes payable $10,000

2.) Account of supplies on December 31,2017, indicates that supplies of $900 are one hand.

Debit Credit
Supplies Expenses $1,550
Supplies $1,550

3.) Depreciation on the equipment for 2017 is $1,000.

Debit Credit
Depreciation Expenses-Equipment $1,000
Accumulated Depreciation-Equipment $1,000

4.) Devin wolf company paid $2,100 for 12 months of insurance coverage on June 1,2017.

2,100 × 7/12=1,225

Debit Credit
Insurance Expense $1,225
Prepaid Expense $1,225

5.) On December 1,2017 Devin wolf company collected$32,000 for consulting service to be performed from December 1,2017,through March 31,2018

32,000/4 months =8,000 per month

Debit Credit
Unearned consulting revenue $8,000
Consulting Fees $8,000

6.) Devin wolf performed consulting services for a client in December 2017.

The client will be billed $4,200

Debit Credit
Accounts receivable $4,200
Consulting fees $4,200

7.) Devin wolf company pays it's employees total salaries of $9,000 every Monday for the preceding 5-day week. On Monday, December 29,employees were paid for the week ending December 26.

All employees worked the last 3 days of 2017.

9,000/5×3 =5,400

Debit Credit
Salaries Expenses $5,400
Salaries Payable $5,400